OpenStudy (anonymous):

Determine whether the graph of the equation is symmetric with respect to the x-axis, the y-axis, and/or the origin. y = x - 5

OpenStudy (anonymous):

 Step 1: Find some points on the line by picking values for one variable, and finding the other: y = x-5 ; y = 0 0 = x - 5 x = 5 (5, 0) y = x - 5 ; x = 0 y = 0 - 5 y = -5 (0, -5) Step 2: Draw it! Quick tips for future: - Symmetric about the x-axis (for a line) means that the equation of the line is of the form \(y=k\) - Symmetric about the y-axis (for a line) means that the equation of the line is of the form \(x=k\) - Symmetric about the origin (for a line) means that the equation of the line is of the form \(y=kx\)